# Thomassen’s proof and Filippov’s proof of the Weak Jordan Theorem

### 05 Apr 2026 v1

@@ Abstract
We present, in detail and with rigour, the two title proofs. The Weak Jordan Theorem states that the complement of any topological circuit in the plane is disconnected.
@@
We denote by $I$ any compact real interval $I=[a,b]$ with $a<b$. If $f\colon I\to X$ is a map, we set $f[I]=\{f(x)\colon\;x\in I\}$, $f^{0}=\{f(x)\colon\;a<x<b\}$ and $\mathrm{ep}(f)=\{f(a),f(b)\}$. We also say that $f$ joins $f(a)$ to $f(b)$ in $X$. If $g\colon A\to B$ is any map and $C$ is any set, we define
@@
An arc is a continuous injection $f\colon I\to\mathbb{R}^{2}$. A circuit is a continuous map $f\colon I\to\mathbb{R}^{2}$ such that $f(x)=f(y)$ for $x\neq y$ if and only if $\{x,y\}=\{a,b\}$. The following Jordan Theorem, or JT, is a fundamental result in plane topology due to C. Jordan [7] in 1893.
Theorem 1.1 (Jordan, 1893)
@@
— it is a disjoint union of two nonempty open connected sets $A$ and $B$, where $A$ is bounded and $B$ is unbounded, such that $\partial A=\partial B=f[I]$.
We explain notions appearing in the statement. We work in the Euclidean plane, which is a metric space $\langle\mathbb{R}^{2},d_{2}\rangle$ with the metric
@@
$X\subset A\cup B\wedge A\cap X\neq\emptyset\neq B\cap X\wedge A\cap B\cap X=\emptyset$
@@
In this article, we will not prove the JT, but we will provide two proofs for the weaker version, the Weak Jordan Theorem, or WJT.
@@
It is clear that JT $\Rightarrow$ WJT because the union of two nonempty open sets that are disjoint is a disconnected set. The proof of the JT is harder than that of the WJT because it requires the following “lemma”, the Arc Theorem, or AT.
@@
We hope to write about this overlooked cornerstone of the theory of planar graphs in a separate contribution [8].
In 1992, C. Thomassen found in [9] an interesting proof of the JT based on planar graphs; he actually obtained a more general result, the Jordan–Schönflies theorem, and the theorem on the classification of surfaces. In Section 2 we present only a part of [9], the proof of the WJT. It can be divided into four claims and employs the following configurations. A $K_{3,3}$-configuration has two three-element sets $U,V\subset\mathbb{R}^{2}$ that are disjoint, and nine arcs $f_{u,v}\colon I\to\mathbb{R}^{2}$ for $u\in U$ and $v\in V$ such that $\mathrm{ep}(f_{u,v})=\{u,v\}$ and
@@
for every two pairs $\langle u,v\rangle\neq\langle u^{\prime},v^{\prime}\rangle$. It is a plane drawing, without crossings, of the complete bipartite graph $K_{3,3}$; these configurations do not exist. A PL map is a map $f\colon[a,b]\to\mathbb{R}^{2}$ such that for a partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $[a,b]$, all restrictions $f\,|\,[a_{i-1},a_{i}]$ for $i=1,2,\dots,k$ are non-constant linear maps. We call $f$ closed if $f(a)=f(b)$. Every PL map is continuous. The straight segments $f[\,[a_{i-1},a_{i}]\,]$ ($\subset\mathbb{R}^{2}$) are called segments of $f[I]$. The points $f(a_{i})$ ($\in\mathbb{R}^{2}$) are called corners of $f[I]$. In a PL $K_{3,3}$-configuration, the nine arcs are PL maps. Thomassen’s proof of the WJT in [9] is the sequence claim 1 $\Rightarrow$ claim 2 $\Rightarrow$ claim 3 $\Rightarrow$ claim 4, where the four claims are as follows.
@@
Another proof of the WJT is contained in the proof of the JT given by A. F. Filippov in [3] in 1950. We will explain this proof in Section 3. It can be divided in three steps.
@@
    such that $N(c)$ is the parity of the number of simple intersections of the vertical half-line going up from the point $c$ ($\in g[I]$) with $f[I]$ (to be defined precisely later). Then $N$ is constant.
@@
    The function $N$ is additive in the PL maps $f$. If $f$ is not closed and $c$ lies between the two vertical lines and below $f[I]$, then $N(c)=1$.
@@
    Let $f\colon I\to\mathbb{R}^{2}$ be a circuit. The results in steps 1 and  2 are used to obtain two distinct points $c,d\in\mathbb{R}^{2}\setminus f[I]$ such that every PL map joining them intersects $f[I]$. This means that the complement $\mathbb{R}^{2}\setminus f[I]$ is disconnected and the WJT holds.
Auxiliary results. In the rest of Section 1 we present, with proofs, several auxiliary results that will be needed in Sections 2 and 3. They are standard, and the impatient reader may now jump to the two proofs and check any auxiliary result if the need arises. We explain the inclusion of the auxiliary results in our article in Section 4.
We begin with three constructions related to maps. Let $I=[a,b]$ and $J=[c,d]$ be real intervals with $a<b$ and $c<d$, and let $f\colon I\to X$ be a map. Then we can parametrize $f[I]$ by $J$ via the map $g=f(h)\colon J\to X$, where $h\colon J\to I$ is the linear homeomorphism $h(x)=\frac{b-a}{d-c}(x-c)+a$. Clearly, $g$ joins $g(c)=f(a)$ to $g(d)=f(b)$. If $f$ is a PL map, then so is $g$, and the same holds for being a continuous map, an arc or a circuit. Thus we can use any compact interval $I=[\alpha,\beta]$, $\alpha<\beta$, as the definition domain of arcs and circuits.
Let $f,g\colon I=[a,b]\to X$ be two maps such that $f(b)=g(a)$. We write $f+g$ for their concatenation. It is the map $h=f+g\colon I\to X$ given by h(x)=f(2xa)h(x)=f(2x-a) for $a\leq x\leq\frac{1}{2}(a+b)$, and by h(x)=g(2xb)h(x)=g(2x-b) for $\frac{1}{2}(a+b)\leq x\leq b$. Clearly, $h$ joins $h(a)=f(a)$ to $h(b)=g(b)$. If $f$ and $g$ are PL maps, then so is $h$, and the same holds for continuity. Let in addition $f[I]\cap g[\,(a,b]\,]=\emptyset$, and let $f$ and $g$ be arcs. Then $f+g$ is an arc. Let in addition $f[I]\cap g^{0}=\emptyset$, $g(b)=f(a)$, and let $f$ and $g$ be arcs. Then $f+g$ is a circuit.
Finally, let $U$, $V$, and $\{f_{u,v}\colon\;u\in V,v\in V\}$ be a $K_{3,3}$-configuration. For any arc $f_{u,v}\colon I=[a,b]\to\mathbb{R}^{2}$ in it, we denote by $f_{v,u}\colon I\to\mathbb{R}^{2}$ the reversed arc
@@
The first group of auxiliary results concerns compactness in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is compact if every sequence in $X$ has a convergent subsequence with a limit in $X$. We take for granted the compactness of $I=[a,b]$ ($\subset\mathbb{R}^{2}$).
@@
Proof. We show that if $X$ is not closed or not bounded, then it is not compact. If $X$ is not closed, then there is a sequence $(a_{n})\subset X$ such that $\lim a_{n}=a\in M\setminus X$. Then $(a_{n})$ does not have any subsequence with a limit in $X$. If $X$ is not bounded, then there exists a sequence $(a_{n})\subset X$ such that $d(a_{m},a_{n})\geq 1$ for every $m<n$. Then $(a_{n})$ does not have any convergent subsequence. $\Box$
@@
Let $\langle M,d\rangle$ be a metric space, $X\subset M$ be a closed set, and $f\colon I=[a,b]\to M$ be a continuous map. If $f[I]\cap X\neq\emptyset$, then the subset $f^{-1}[X]$ of $I$ has both minimum and maximum.
@@
We omit the proof because we prove the following more general statement.
@@
In Thomassen’s proof, Proposition 1.8 suffices. In Filippov’s proof, we need Proposition 1.9.
Recall that a map $f\colon M\to N$ between metric spaces $\langle M,d\rangle$ and $\langle N,e\rangle$ is continuous at a point $a\in M$ if for every $\varepsilon>0$ there is $\delta>0$ such that $f[B(a,\delta)]\subset B(f(a),\varepsilon)$. The map $f$ is continuous if it is continuous at every point in $M$. The map $f$ is uniformly continuous if for every $\varepsilon>0$ there is $\delta>0$ such that $d(x,y)<\delta$ $\Rightarrow$ $e(f(x),f(y))<\varepsilon$ for every $x,y\in M$. It is easy to show that for every continuous map $f$ and every open (closed) set $X\subset N$ the preimage $f^{-1}[X]$ is an open (closed) set in $M$. Also, if $f$ is continuous at $a$ then for every sequence $(a_{n})\subset M$ with $\lim a_{n}=a$ we have $\lim f(a_{n})=f(a)$.
@@
Proof. Suppose, for the contrary, that $f$ is not uniformly continuous. Then there exists $\varepsilon>0$ and sequences $(a_{n}),(b_{n})\subset M$ such that $\lim d(a_{n},b_{n})=0$, but $e(f(a_{n}),f(b_{n}))\geq\varepsilon$ for every $n$. Due to the compactness of $M$ we may assume that $\lim a_{n}=\lim b_{n}=a$ ($\in M$). But $\lim f(a_{n})=\lim f(b_{n})=f(a)$ ($\in N$) does not hold, in contradiction with the continuity of $f$ at $a$ $\Box$
The second group of auxiliary results concerns connectedness.
@@
As an application, we prove the connectedness of balls in the plane.
@@
Proof. Let $B=B(b,r)=\{a\in\mathbb{R}^{2}\colon\;d_{2}(a,b)<r\}$ be a ball in the Euclidean plane. For any point $a\in B$ we denote by $S_{a}$ the (possibly degenerate) straight segment joining $b$ and $a$. Every $S_{a}$ is connected by Proposition 1.11, as a continuous image of an interval. Thus
$B=\bigcup_{a\in B}S_{a}$}
@@
Let $X\subset\mathbb{R}^{2}$ be an open connected set and let $u,v\in X$ with $u\neq v$. Then a PL arc joins $u$ to $v$ in $X$.
Proof. Let $Y\subset X$ be the set of $x\in X$ such that $x=u$ or a PL arc joins $u$ to $x$ in $X$. Clearly, $Y\neq\emptyset$ and $Y$ is open. We show that also the set $Z=X\setminus Y$ is open. If $Z\neq\emptyset$, we have the contradiction that $X$ is disconnected. Thus $Y=X$ and a PL arc joins $u$ to $v$ in $X$.
Let $y\in Z$. We take a ball $B=B(y,r)\subset X$. Let $z\in B$ with $z\neq y$. Suppose that $f\colon I\to X$ is a PL arc joining $u$ to $z$ in $X$. Let $s$ ($\subset B$) be the shortest straight segment joining $y$ and $f[I]$. Using $f$ and $s$ we could obtain a PL arc joining $u$ to $y$ in $X$, which contradicts $y\in Z$. Thus $f$ does not exist and $z\in Z$. We see that $B\subset Z$ and $Z$ is open. $\Box$
The last group of auxiliary results is related to Thomassen’s construction of a plane drawing of $K_{3,3}$ from a circuit whose complement is connected.
@@
According to the initial remark, we may redefine $g$ and $h$ on the domain $I$. $\Box$
@@
Let $f\colon I=[a,b]\to\mathbb{R}^{2}$ be continuous, $f(a)_{y}<f(b)_{y}$ and let $\ell$ be a horizontal line with the $y$-coordinate in $(f(a)_{y},f(b)_{y})$. Then $\ell\cap f[I]\neq\emptyset$.
Proof. Apply to the continuous function $g\colon I\to\mathbb{R}$, where $g(t)=f(t)_{y}$, the intermediate value theorem. $\Box$
@@
Let $A,B\subset[0,1]$ be nonempty closed sets that are disjoint. Then there exists a subinterval $x,y]\subset[0,1]$ such that $x\in A$ and $y\in B$, or $x\in B$ and $y\in A$, and $(x,y)\cap(A\cup B)=\emptyset$.

### 19 Apr 2026 v3

@@ Abstract
We present, in detail and with a modern rigor, the two title proofs. The Weak Jordan Theorem (WJT) states that the complement of any topological circuit in the plane is disconnected.
@@
Let us begin with a few necessary definitions. We denote by $I$ any compact real interval $I=[a,b]$ with $a<b$. If $f\colon I\to X$ is a map, we set $f[I]=\{f(x)\colon\;x\in I\}$, $f^{0}=\{f(x)\colon\;a<x<b\}$ and $\mathrm{ep}(f)=\{f(a),f(b)\}$. We also say that $f$ joins $f(a)$ to $f(b)$ in $X$. If $f(a)=f(b)$, we call $f$ closed. If $g\colon A\to B$ is any map and $C$ is any set, we define
@@
An arc is a continuous injection $f\colon I\to\mathbb{R}^{2}$. A circuit is a continuous map $f\colon I=[a,b]\to\mathbb{R}^{2}$ such that $f(x)=f(y)$ for $x\neq y$ if and only if $\{x,y\}=\{a,b\}$. The following Jordan Theorem, or JT, is a fundamental result in plane topology. It is due to C. Jordan [14, pp. 587–594] in 1887.
Theorem 1.1 (Jordan, 1887)
@@
is a disjoint union of two nonempty open connected sets $A$ and $B$ such that $A$ is bounded, $B$ is unbounded, and $\partial A=\partial B=f[I]$.
We explain the statement. We work in the Euclidean plane, which is a metric space $\langle\mathbb{R}^{2},d_{2}\rangle$ with the metric
@@
$A\cup B\supset X\;\&\;A\cap X\neq\emptyset\neq B\cap X\;\&\;A\cap B\cap X=\emptyset$
@@
In this article, we will not prove the JT, but we present two proofs for the weaker version, the Weak Jordan Theorem, or WJT.
@@
JT $\Rightarrow$ WJT because the disjoint union of two nonempty open sets is disconnected. The proof of the JT is harder than that of the WJT because it requires the following “lemma”, the Arc Theorem, or AT.
@@
We write about this overlooked cornerstone of the theory of planar graphs in a separate contribution [15], where we also complete the presentation of a proof of the JT.
In 1992, C. Thomassen found in [17] an interesting proof of the JT based on planar graphs. He actually obtained combinatorial proofs of the more general Jordan–Schönflies theorem and of the theorem on classification of surfaces. In Section 2 we present only a part of [17], Thomassen’s proof of the WJT. It employs the following virtual configurations. A $K_{3,3}$-configuration has two disjoint three-element sets $U,V\subset\mathbb{R}^{2}$ and nine arcs $f_{u,v}\colon I\to\mathbb{R}^{2}$ for $u\in U$ and $v\in V$ such that $f_{u,v}$ joins $u$ to $v$ and
@@
for every two pairs $\langle u,v\rangle\neq\langle u^{\prime},v^{\prime}\rangle$. It is a plane drawing, without crossings, of the complete bipartite graph $K_{3,3}$; these configurations do not exist. A PL map is a map $f\colon[a,b]\to\mathbb{R}^{2}$ such that for a partition $a=a_{0}<a_{1}<\dots<a_{k}=b$ of $[a,b]$, all restrictions $f\,|\,[a_{i-1},a_{i}]$ for $i=1,2,\dots,k$ are non-constant linear maps. Every PL map is continuous. The straight segments $f[\,[a_{i-1},a_{i}]\,]$ ($\subset\mathbb{R}^{2}$) are the segments of $f[I]$. The points $f(a_{i})$ ($\in\mathbb{R}^{2}$) are the corners of $f[I]$. In a PL $K_{3,3}$-configuration, the nine arcs are PL maps. Thomassen’s proof of the WJT in [17], presented in Section 2, is the sequence claim 1 $\Rightarrow$ claim 2 $\Rightarrow$ claim 3 $\Rightarrow$ claim 4, where the four claims are as follows.
@@
Another proof of the WJT is contained in the proof of the JT given by A. F. Filippov in [5] in 1950. We present this proof in Section 3. It can be divided in three steps.
@@
    such that $N(c)$ is the parity of the number of simple intersections of the vertical half-line going up from the point $c$ with $f[I]$ (we precisely define $N$ later). Then $N$ is constant.
@@
    The function $N(c,f)$ is additive in the variable $f$. If $f$ is not closed and the point $c$ lies between the two vertical lines and below all intersections of the line $x=c_{x}$ with $f[I]$, then $N(c)=1$.
@@
    Let $f\colon I\to\mathbb{R}^{2}$ be a circuit. The results in steps 1 and 2 are used to obtain two distinct points $c,d\in\mathbb{R}^{2}\setminus f[I]$ such that for every PL map $g$ joining $c$ to $d$ we have $g[I]\cap f[I]\neq\emptyset$. This means that the complement $\mathbb{R}^{2}\setminus f[I]$ is disconnected and the WJT holds.
Why this presentation of the two proofs? They are beautiful, but they are not rigorous by modern standards. Since we like them, we decided to elevate them to the present level of rigor. See also the discussion in Section 4.
Auxiliary results. In the rest of Section 1 we present, with proofs, auxiliary results that will be needed in Sections 2 and 3. Since they are mostly standard, the impatient reader may now jump to the two proofs and check any auxiliary result only if the need arises.
We begin with three constructions related to maps. Let $I=[a,b]$ and $J=[c,d]$ be real intervals with $a<b$ and $c<d$, and let $f\colon I\to X$ be a map. We can parametrize $f[I]$ by $J$ via the map $g=f(h)\colon J\to X$, where $h\colon J\to I$ is the linear homeomorphism $h(x)=\frac{b-a}{d-c}(x-c)+a$. Clearly, $g$ joins $g(c)=f(a)$ to $g(d)=f(b)$ in $X$. If $f$ is a PL map, then so is $g$, and the same holds for being a continuous map, an arc, or a circuit. Thus we can use any compact interval $I=[\alpha,\beta]$, $\alpha<\beta$, as the common definition domain of PL maps, (PL) arcs, and (PL) circuits.
Let $f,g\colon I=[a,b]\to X$ be two maps such that $f(b)=g(a)$. We write $f+g$ for their concatenation $h=f+g\colon I\to X$ given by h(x)=f(2xa)h(x)=f(2x-a) for $a\leq x\leq\frac{1}{2}(a+b)$, and by h(x)=g(2xb)h(x)=g(2x-b) for $\frac{1}{2}(a+b)\leq x\leq b$. Clearly, $h$ joins $h(a)=f(a)$ to $h(b)=g(b)$ in $X$. If $f$ and $g$ are PL maps, then so is $h$, and the same holds for continuity. Let in addition $f[I]\cap g[\,(a,b]\,]=\emptyset$, and let $f$ and $g$ be arcs. Then $f+g$ is an arc. Let in addition $f[I]\cap g^{0}=\emptyset$, $g(b)=f(a)$, and let $f$ and $g$ be arcs. Then $f+g$ is a circuit.
Finally, let $U$, $V$, and $\{f_{u,v}\colon\;u\in V,v\in V\}$ be a $K_{3,3}$-configuration. For any of the arcs $f_{u,v}\colon I=[a,b]\to\mathbb{R}^{2}$, we denote by $f_{v,u}\colon I\to\mathbb{R}^{2}$ the reverse arc
@@
The first two auxiliary results concern closed (and non-compact) sets in metric spaces used in Filippov’s proof, such as vertical half-lines in $\mathbb{R}^{2}$. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is closed if the complement $M\setminus X$ is an open set, that is, if for every point $a\in M\setminus X$ there exists an $r>0$ such that the ball $B(a,r)\subset X\setminus M$.
Proposition 1.4
In any metric space, the union of two closed sets is a closed set.
Proof. Let $X,Y\subset M$ be closed sets in a metric space $\langle M,d\rangle$, and let
$a\in M\setminus(X\cup Y)=(M\setminus X)\cap(M\setminus Y)\,.$
We have $B(a,r)\subset M\setminus X$ and $B(a,r^{\prime})\subset M\setminus Y$ for some $r,r^{\prime}>0$. Thus
$B(a,\,r^{\prime\prime})\subset(M\setminus X)\cap(M\setminus Y)$
for $r^{\prime\prime}=\min(r,r^{\prime})$ (>0>0). We see that $X\cup Y$ is closed. $\Box$
Proposition 1.5
Let $\ell\subset\mathbb{R}^{2}$ be a vertical half-line going up or down from a point $c\in\mathbb{R}^{2}$. Then $\ell$ is a closed set.
Proof. Suppose that $\ell$ goes down (if it goes up, we argue similarly) and that $a\in\mathbb{R}^{2}\setminus\ell$. If $a_{x}=c_{x}$, we define $r=a_{y}-c_{y}$ (>0>0), and if $a_{x}\neq c_{x}$, we set $r=|a_{x}-c_{x}|$ (>0>0). Then $d_{2}(a,b)\geq r$ for every point $b\in\ell$, and therefore $B(a,r)\subset\mathbb{R}^{2}\setminus\ell$. $\Box$
The second group of auxiliary results concerns compactness in metric spaces. Recall that a set $X\subset M$ in a metric space $\langle M,d\rangle$ is compact if every sequence in $X$ has a convergent subsequence with a limit in $X$. We take for granted the compactness of $I=[a,b]$ ($\subset\mathbb{R}^{2}$).
@@
Proof. We show that if $X$ is not closed or not bounded, then it is not compact. If $X$ is not closed, then there is a sequence $(a_{n})\subset X$ such that $\lim a_{n}=a\in M\setminus X$. Such $(a_{n})$ does not have any subsequence with a limit in $X$. If $X$ is not bounded, then there exists a sequence $(a_{n})\subset X$ such that $d(a_{m},a_{n})\geq 1$ for every $m<n$. Such $(a_{n})$ does not have any convergent subsequence. $\Box$
@@
Let $\langle M,d\rangle$ be a metric space, $X\subset M$ be a closed set, and let $f\colon I=[a,b]\to M$ be a continuous map. If $f[I]\cap X\neq\emptyset$, then the subset $f^{-1}[X]$ of $I$ has both minimum and maximum.
@@
We omit the proof because we prove a stronger result.
@@
In Thomassen’s proof, Proposition 1.10 suffices. Filippov’s proof uses also Proposition 1.11.
Recall that a map $f\colon M\to N$ between metric spaces $\langle M,d\rangle$ and $\langle N,e\rangle$ is continuous at a point $a\in M$ if for every $\varepsilon>0$ there is $\delta>0$ such that $f[B(a,\delta)]\subset B(f(a),\varepsilon)$. The map $f$ is continuous if it is continuous at every point in $M$. The map $f$ is uniformly continuous if for every $\varepsilon>0$ there is $\delta>0$ such that $d(x,y)\leq\delta$ $\Rightarrow$ $e(f(x),f(y))\leq\varepsilon$ for every $x,y\in M$. It is easy to show that for every continuous map $f$ and every open (closed) set $X\subset N$ the preimage $f^{-1}[X]$ is an open (closed) set in $M$. If $f$ is continuous at $a$ then for every sequence $(a_{n})\subset M$ with $\lim a_{n}=a$ we have $\lim f(a_{n})=f(a)$.
@@
Proof. Suppose, for the contrary, that $f$ is not uniformly continuous. Then there exist $\varepsilon>0$ and sequences $(a_{n}),(b_{n})\subset M$ such that $\lim d(a_{n},b_{n})=0$, but $e(f(a_{n}),f(b_{n}))\geq\varepsilon$ for every $n$. Due to the compactness of $M$ we may assume that $\lim a_{n}=\lim b_{n}=a$ ($\in M$). But $\lim f(a_{n})=\lim f(b_{n})=f(a)$ ($\in N$) does not hold, in contradiction with the continuity of $f$ at $a$ $\Box$
The third group of auxiliary results concerns connectedness.
@@
We prove, as an application, that balls in the plane are connected.
@@
Proof. Let $B=B(b,r)=\{a\in\mathbb{R}^{2}\colon\;d_{2}(a,b)<r\}$ be a ball in the Euclidean plane. For any point $a\in B\setminus\{b\}$ we denote by $S_{a}$ the straight segment joining $b$ and $a$. Every $S_{a}$ is connected by Proposition 1.13, as a continuous image of an interval. Thus
$B=\bigcup_{a\in B\setminus\{b\}}S_{a}$}
@@
The following proposition is the most intricate of the auxiliary results. We need a lemma for it.
Lemma 1.17
Let $X\subset\mathbb{R}^{2}$ be an open set, let $u,v\in X$ with $u\neq v$, and let there be a PL arc joining $u$ to $v$ in $X$. Then there exists a $v$-centered ball $B\subset X$ such that for every point $w\in B$ there exists a PL arc joining $u$ to $w$ in $X$.
Proof. Let $f\colon I=[a,b]\to X$ be a PL arc joining $u$ to $v$ in $X$, let $a=a_{0}<a_{1}<\dots<a_{k}=b$ be the partition of $[a,b]$, and let $s_{i}=f[\,[a_{i-1},a_{i}]\,]$ for $i=1,2,\dots,k$ be the segments of $f[I]$. Since $d_{2}(s_{i},v)>0$ for $i=1,2,\dots,k-1$ (Propositions 1.6 and 1.10), we can take small enough $r>0$ such that $B=B(v,r)\subset X$ and $s_{i}\cap B=\emptyset$ for $i=1,2,\dots,k-1$. We define for every point $w\in B$ a PL arc $f_{w}$ joining $u$ to $w$ in $X$. If $w\in B\cap s_{k}$, then $f_{w}$ is the initial part of $f$ ending at $w$. If $w\in B\setminus s_{k}$, then $f_{w}$ is the extension of $f$ by the straight segment joining $v$ to $w$. It is easy to see that $f_{w}$ is as claimed. $\Box$
@@
Let $X\subset\mathbb{R}^{2}$ be an open connected set and let $u,v\in X$ with $u\neq v$. Then there exists a PL arc joining $u$ to $v$ in $X$.
Proof. Let $Y\subset X$ be the set of points $x\in X$ such that $x=u$ or a PL arc joins $u$ to $x$ in $X$. Clearly, $Y\neq\emptyset$ ($u\in Y$) and $Y$ is open (Lemma 1.17). We show that also the set $Z=X\setminus Y$ is open. If $Z\neq\emptyset$, we have the contradiction that $X$ is disconnected. Thus $Z=\emptyset$ and $Y=X$, which implies that there exists a PL arc joining $u$ to $v$ in $X$.
Let $y\in Z$. We take a ball $B=B(y,r)\subset X$ such that $u\not\in B$, which is possible as $y\neq u$, and show that $B\subset Z$. Let $z\in B$ with $z\neq y$. Clearly, $z\neq u$. Suppose that $f\colon I\to X$ is a PL arc joining $u$ to $z$ in $X$. Let $s=yz$ ($\subset B$) be the straight segment joining $y$ and $z$. Considering the first intersection in $f[I]\cap s$ (Proposition 1.9), we obtain a PL arc joining $u$ to $y$ in $X$, but this contradicts $y\in Z$. Thus $f$ does not exist and $z\in Z$. We see that $B\subset Z$. The set $Z$ is open. $\Box$
The last group of auxiliary results is related to Thomassen’s construction of a plane drawing of $K_{3,3}$ based on a circuit whose complement is connected.
Proposition 1.19
Let $X\subset\mathbb{R}^{2}$ be a compact set. Then there exist four points $u,v,r,s\in X$ such that $u_{x}$ and $v_{x}$ have, respectively, the maximum and minimum $x$-coordinate on $X$, and $r$ and $s$ have an analogous property in the $y$-coordinate.
Proof. We prove the existence of $v$ (the existence of the other three points is established similarly). Since $X$ is bounded (Proposition 1.7), there exists a sequence $(a_{n})\subset X$ such that
$\lim_{n\to\infty}(a_{n})_{x}=\inf(\{a_{x}\colon\;a\in X\})=:\alpha\,.$
Since $X$ is compact, we may assume that $\lim a_{n}=v\in X$. Since the function $c\mapsto c_{x}$ is continuous,
$v_{x}=\big(\lim_{n\to\infty}a_{n}\big)_{x}=\lim_{n\to\infty}(a_{n})_{x}=\alpha$
and $v$ is a leftmost point of $X$. $\Box$
@@
According to the initial remark, we may redefine $g$ and $h$ on $I$. It is easy to see that $g$ and $h$ are as stated. $\Box$
@@
Let $f\colon I=[a,b]\to\mathbb{R}^{2}$ be continuous, $f(a)_{y}<f(b)_{y}$, and let $\ell$ be a horizontal line with the $y$-coordinate in $(f(a)_{y},f(b)_{y})$. Then $\ell\cap f[I]\neq\emptyset$.
Proof. We apply the intermediate value theorem to the continuous function $f(t)_{y}\colon I\to\mathbb{R}$. $\Box$
@@
Let $A,B\subset[c,d]$ be nonempty closed sets that are disjoint. Then there exists a subinterval $x,y]\subset[c,d]$ such that $x\in A$ and $y\in B$, or $x\in B$ and $y\in A$, and $(x,y)\cap(A\cup B)=\emptyset$.